For a particle in one-dimensional box, write the equations for calculating the energy values corresponding to i) n = 1 ii) n = 2 iii) n = 3
1
Expert's answer
2020-02-11T05:38:08-0500

En= (n^2 x h^2)/(8 x m x L^2)

E1= (1^2 x h^2)/(8mL^2)

E2=(2^2 x h^2)/(8mL^2)

E3=(3^2 x h^2)/(8mL^2)